# An infinite family of pairs of distinct quartic Galois CM-fields with the same discriminant and regulator

### v1

@@ 6. The regulator of the real quadratic field $\mathbb{Q}(\sqrt{t^{2}+1})$
$\mathrm{reg}\left(\mathbb{Q}(\sqrt{t^{2}+1})\right)=\log(t+\sqrt{t^{2}+1})>\log t>M.$
@@ 7. Proofs of the main theorems and examples
Let $M$ be a positive real number. By Proposition 6.4, there is a positive integer $t$ such that $t\equiv 5\pmod{8}$ and the regulator $R$ of the real quadratic field $\mathbb{Q}(\sqrt{t^{2}+1})$ is greater than $M$. Furthermore, by Dirichlet’s theorem on arithmetic progressions, there are infinitely many primes $p$ satisfying $p>t^{2}+1$ (hence in particular, $\gcd(p,t^{2}+1)=1$) and $p\equiv 1\pmod{4}$. For each such prime $p$, the biquadratic fields $B_{-p,t^{2}+1}$ are distinct because different values of $p$ yield different combinations of three quadratic subfields. By the discussion in Section 3, these biquadratic fields have the same maximal real subfield $\mathbb{Q}(\sqrt{t^{2}+1})$ and hence the same regulator $R$. This proves (i).
@@
Let $M$ be a positive real number. By Proposition 6.4, there is a positive integer $t$ such that $t\equiv 5\pmod{8}$ and the regulator $R$ of the real quadratic field $\mathbb{Q}(\sqrt{t^{2}+1})$ is greater than $M$. For each odd prime $p$ satisfying $p>t^{2}+1$, the cyclic quartic fields $K_{-p,t}$ are distinct because their discriminants are distinct by Proposition 5.2. It follows from Lemma 4.2 that all such fields $K_{-p,t}$ have the same maximal real subfield $\mathbb{Q}(\sqrt{t^{2}+1})$. Since all of them are CM-fields, by Theorem 2.1, they have the same regulator $R$. This proves (i).
@@
We consider the pair of the cyclic quartic fields $K_{-3,35}$ and $K_{-6,35}$. They share the maximal real subfield $\mathbb{Q}(\sqrt{2\cdot 613})$, whose fundamental unit is $\varepsilon=35+\sqrt{2\cdot 613}$. Moreover, they have the same discriminant $2^{11}\cdot 3^{2}\cdot 13^{3}$, regulator $2\cdot\log\lvert\varepsilon\rvert=8.4973985\cdots$ and class number $2^{3}\cdot 5^{2}\cdot 97$.
@@
For an odd integer $t$ greater than $1$, the discriminant of the real quadratic field $\mathbb{Q}(\sqrt{t^{2}+1})$ has more than one prime factor, hence by genus theory, its class number is even.
@@ Acknowledgements
The authors would like to thank Professor Shin Nakano for his valuable comments that improved the manuscript.

### v2

@@ 6. The regulator of the real quadratic field $\mathbb{Q}(\sqrt{t^{2}+1})$
This classical result can be found in Degert [4, p. 6] or Hasse [6, p. 49]; we include a proof for the reader’s convenience.
@@
$\operatorname{reg}\left(\mathbb{Q}(\sqrt{t^{2}+1})\right)=\log(t+\sqrt{t^{2}+1})>\log t>M.$
Remark 6.5.
An analogous result holds if we replace the condition $t\equiv 5\pmod{8}$ with $t\equiv 3\pmod{8}$ in Proposition 6.4. This follows by replacing Lemma 6.2 with its analogue for $t\equiv 3\pmod{8}$.
@@ 7. Proofs of the main theorems and examples
Let $M$ be a positive real number. By Proposition 6.4, there is a positive integer $t$ such that $t\equiv 5\pmod{8}$ and the regulator $R_{0}$ of the real quadratic field $\mathbb{Q}(\sqrt{t^{2}+1})$ is greater than $M$. Furthermore, by Dirichlet’s theorem on arithmetic progressions, there are infinitely many primes $p$ satisfying $p>t^{2}+1$ (hence in particular, $\gcd(p,t^{2}+1)=1$) and $p\equiv 1\pmod{4}$. For each such prime $p$, the biquadratic fields $B_{-p,t^{2}+1}$ are distinct because different values of $p$ yield different combinations of three quadratic subfields. By the discussion in Section 3, these biquadratic fields have the same maximal real subfield $\mathbb{Q}(\sqrt{t^{2}+1})$ and hence the same regulator $R=2R_{0}$. This proves (i).
@@
Let $M$ be a positive real number. By Proposition 6.4, there is a positive integer $t$ such that $t\equiv 5\pmod{8}$ and the regulator $R_{0}$ of the real quadratic field $\mathbb{Q}(\sqrt{t^{2}+1})$ is greater than $M$. For each odd prime $p$ satisfying $p>t^{2}+1$, the cyclic quartic fields $K_{-p,t}$ are distinct because their discriminants are distinct by Proposition 5.2. It follows from Lemma 4.2 that all such fields $K_{-p,t}$ have the same maximal real subfield $\mathbb{Q}(\sqrt{t^{2}+1})$. Since all of them are CM-fields, by Theorem 2.1, they have the same regulator $R=2R_{0}$. This proves (i).
@@
We consider the pair of the cyclic quartic fields $K_{-3,35}$ and $K_{-6,35}$. They share the maximal real subfield $\mathbb{Q}(\sqrt{2\cdot 613})$, whose fundamental unit is $\varepsilon=35+\sqrt{2\cdot 613}$. Moreover, they have the same discriminant $2^{11}\cdot 3^{2}\cdot 613^{3}$, regulator $2\cdot\log\lvert\varepsilon\rvert=8.4973985\cdots$ and class number $2^{3}\cdot 5^{2}\cdot 97$.
@@
For an odd integer $t$ greater than $1$, if $t^{2}+1\not\in 2\mathbb{Q}^{2}$, the discriminant of the real quadratic field $\mathbb{Q}(\sqrt{t^{2}+1})$ has more than one prime factor, and hence by genus theory the narrow class number is even. Since the fundamental unit $t+\sqrt{t^{2}+1}$ of $\mathbb{Q}(\sqrt{t^{2}+1})$ has norm 1-1, the narrow and usual class numbers coincide, and therefore the class number is even.
@@ Acknowledgements
The authors would like to thank Professor Shin Nakano for his valuable comments and the anonymous referee for careful reading and helpful suggestions.